OpenStudy (mtalhahassan2):

Explain what is meant by “field” and compare the properties of gravitational, electric, and magnetic forces in terms of particles affected, factors affecting the magnitude, and their relative strengths.

OpenStudy (mtalhahassan2):

@nuttyliaczar

OpenStudy (emyoung0):

gravitational and electric fields are similar. They are both inversely proportional to the square of the distance, and directly proportional to the mass or charge. G = kM/r^2 E = k*Q/d^2

OpenStudy (mtalhahassan2):

LOL I know that

OpenStudy (mtalhahassan2):

do you know another information?

OpenStudy (zzr0ck3r):

I field is a commutative ring such that there exists a unit and everything has a multiplicative inverse.

OpenStudy (thomas5267):

Not that field zzr0ck3r =_=

OpenStudy (zzr0ck3r):

oh I just figured because this is for math...you know... :)